# Concept-based Adversarial Attack: a Probabilistic Perspective

### v1

@@ 4.1 Augment Concept Datasets by Modern Generative Models
In practice, it can be somewhat challenging to obtain a high-quality, highly diverse dataset $\mathcal{C}_{\text{ori}}$ depicting the same concept, as required by our method. For example, as shown on the left side of Figure 2, the dataset provided by DreamBooth [32] contains five images of the same long-eared corgi. Although the corgi is shown in various poses and from multiple viewpoints, the relatively uniform backgrounds do not provide sufficient diversity for our concept-based adversarial attack. Therefore, we decided to use Stable Diffusion XL [29] to expand the concept-description dataset.
As illustrated in Figure 2, we designate the corgi as “[V] dog”. Using LoRA finetuning [16], we train an SDXL LoRA model on this concept. Next, we feed the five corgi images into GPT-4o [18], stating that these images represent the “[V] dog” and asking it to produce SDXL prompts that embody sufficient diversity for the “[V] dog.” Finally, we load the corgi LoRA into SDXL and, guided by GPT-4o’s prompts (on the top of Figure 2), generate images featuring a wide range of viewpoints, environments, and poses for this corgi concept (see the right side of Figure 2).
@@ 4.2 Sample Selection
A key advantage of probabilistic adversarial attacks is that we can draw multiple samples from $p_{\text{adv}}$ and select the best ones. In a white-box scenario, we can simply discard samples that fail to deceive the classifier (rejection sampling). However, if $p_{\text{dis}}$ and $p_{\text{vic}}$ overlap only slightly, this may lead to high rejection rates (especially under a top-1 success criterion).
@@ 5.1 Data Preparation
We choose the DreamBooth [32] dataset, which contains 30 different objects to represent the concepts. Each object is initially described by 5 to 6 images; these objects may be animals, dolls, or everyday items such as vases or candles. Because each object has distinctive individual characteristics, the DreamBooth dataset is well-suited for representing concepts. We then apply the augmentation method described in Section 4.1 to generate additional images. For each concept, we produce 30 augmented images, excluding the 5–6 original images to ensure more diverse scenes. We refer to this new, fully augmented dataset as DreamBoothPlus. Among the 30 objects of DreamBooth, we only augment 26, excluding four that pose challenges for text generation or require different fine-tuning parameters for cartoon-style content.
@@ 5.2 Fitting the Distance Distributions
We use the DreamBoothPlus dataset to finetune the diffusion model [10, 28] to fit the distance distribution $p_{\text{dis}}$ (Details in Appendix D). We choose this model over more advanced architectures, such as the Stable Diffusion series [29, 31] or Flux, because it directly models $p(x)$ instead of $p(x\mid y)$, where $x$ is the image and $y$ is a label or prompt. Our goal is to employ a more principled model to illustrate our general adversarial attack method, rather than to optimize for the highest possible engineering performance.
@@ 5.4 Generating Targeted Adversarial Examples
We evaluate the performance of the concept-based adversarial attack in a targeted adversarial attack setting, because targeted attacks are generally more difficult than untargeted attacks [^1] This is especially true for ImageNet classifiers, which must distinguish among 1,000 classes. In an untargeted attack, the goal is simply to prevent the victim classifier from assigning the adversarial sample to its correct class. In contrast, a targeted attack requires the classifier to misclassify the adversarial sample exactly as the chosen target class $y_{\text{tar}}$.. In our experiments, we compare NCF [47], ACA [7], DiffAttack [6], and ProbAttack [49]. We include NCF because it is the strongest color-based adversarial attack. Both ACA and DiffAttack apply adversarial gradients in the latent space induced by Stable Diffusion and DDIM, representing the state of the art in unrestricted adversarial attacks. As discussed earlier in Section 3, ProbAttack can be viewed as a special case of our approach when $|\mathcal{C}_{\text{ori}}|=1$. For both ProbAttack and our approach, we set the number of samples $M$ to 10. During the sample selection phase of the experiments, our method uses two strategies — a conservative strategy and an aggressive strategy (both described in Section 4.2) — denoted in the tables as OURS (CONS) and OURS (AGGR), respectively.
@@
In our experiments, ResNet50 [12] is used as the victim classifier for white-box attacks. We measure transferability on VGG19 [35], ResNet152 [12], DenseNet161 [17], Inception V3 [41], EfficientNet B7 [43], and on adversarially trained Inception V3 Adv [21], EfficientNet B7 Adv [46], and Ensemble IncRes V2 [44]. We report both the white-box targeted attack success rate (on ResNet50) and the black-box transfer success rate (on the remaining models).
Table 1: Targeted attack success rates (%) on ImageNet classifiers. In the white-box setting, a targeted attack is counted as successful if the target class is ranked first. For transferability, we report top-100 success rates, counting an attack as successful if the target class is among the top 100 predictions (since top-1 success was uniformly low across all methods).
@@
Transferability Top100
VGG19 16.41 16.28 31.41 18.97 20.90 34.49
ResNet 152 17.44 21.03 47.69 26.79 35.13 54.10
DenseNet 161 16.15 19.74 43.46 33.08 41.03 63.21
Inception V3 11.41 15.90 24.23 18.21 19.87 25.64
EfficientNet B7 16.41 16.92 22.82 17.44 20.38 28.46
Adversarial Defence Top100
Inception V3 Adv 14.49 16.92 24.74 20.00 20.51 28.33
EfficientNet B7 Adv 11.54 14.10 22.69 21.03 24.74 38.46
Ensemble IncRes V2 12.44 14.62 20.00 17.31 17.44 27.31
For unrestricted adversarial examples, we must also check whether they preserve the original concept and remain undetectable to humans. Therefore, we measure (1) similarity to the original image/concept through a user study (see Appendix F) and CLIP [30] similarity, and (2) image quality using no-reference metrics (MUSIQ [19], TReS [11], NIMA [42], ARNIQA [1], DBCNN [50], and HyperIQA [39]).
As ImageNet has 1,000 classes, it is impractical to evaluate them all. Therefore, we randomly select 30 target classes $y_{\text{tar}}$, listed in Appendix E. Since DreamBoothPlus contains 26 concepts, each method generates $26\times 30=780$ adversarial examples. This scale is comparable to current popular approaches performing untargeted adversarial attacks on the ImageNet-Compatible dataset [22].
Table 1 presents the targeted attack success rates on ImageNet classifiers. Since the choice of aggressive or conservative strategy in the sample selection phase does not affect white-box performance, those rates are identical. Notably, the aggressive strategy achieves significantly higher transferability than other methods. While the conservative strategy leads to slightly lower transferability, it is still roughly comparable to the baseline methods.
@@
$x_{\text{ori}}$ / $\mathcal{C}_{\text{ori}}$$y_{\text{tar}}$iffAtkProbAtkOURS(CONS)OURS(AGGR)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionambulanceRefer to captionbasenjiRefer to captionhamsterRefer to captionking snakeRefer to captionpandaRefer to captiongoldfishRefer to captionmushroomRefer to captionlaptopRefer to captionpapillonRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 3: Qualitative comparison. (A green border indicates an example that successfully fools the classifier; red indicates failure.)
@@ 6 Related Works
To the best of our knowledge, no prior adversarial attack research has defined a distance between an adversarial example and a concept described by a set of images. The most closely related work is [38]. However, Song et al. [38] treats a “class” (e.g., cat, dog, or truck) as the concept, which cannot precisely capture something like an individual identity. In our approach, we represent a concept using a set of images or a probabilistic model, allowing it to be either an identity or a class. By contrast, other unrestricted adversarial attack methods focus solely on generating adversarial examples from a single image. Even advanced methods such as ACA [7] and DiffAttack [6] simply apply the attack gradient to the latent space induced by the DDIM [36] from a single image. Color-based transformations have proven effective in preserving semantic content for untargeted attacks [2, 15, 34, 47, 51], yet they perform poorly in targeted scenarios [6], a result confirmed by our experiments.
Our approach is directly inherited from [49]. The probabilistic perspective of adversarial attacks proposed by [49] serves as a bridge between our concept-based adversarial attack and traditional adversarial attacks. Conceptually, however, our main difference from Zhang et al. is that they aim to leverage prior knowledge from PGM to make semantically meaningful modifications to individual images — an approach essentially similar to ACA and DiffAttack — whereas our work highlights the importance of constructing a concept-based distribution. From this probabilistic viewpoint, we emphasize that adversarial examples arise at the intersection of $p_{\text{dis}}$ and $p_{\text{vic}}$, and we strive to expand this intersection at the semantic level.
7 Limitations
8 Practical Usage Guidelines
@@ Appendix A Proof of the Theorems
Let $p$ be a probability distribution and $q$ be a Gaussian distribution with mean $\mu$ and variance $\sigma^{2}$. Then $KL(p||q)$ is a decreasing function of $\sigma^{2}$ when $\sigma^{2}<\mathbb{E}_{X\sim p}[(X-\mu)^{2}]$.
@@
[…]
$=\mathbb{E}_{X\sim p}\left[\frac{1}{2}\log(2\pi\sigma^{2})+\frac{(X-\mu)^{2}}{2\sigma^{2}}\right]$
$=\frac{1}{2}\log(2\pi\sigma^{2})+\frac{\mathbb{E}_{X\sim p}\left[(X-\mu)^{2}\right]}{2\sigma^{2}}$
Taking the derivative with respect to $\sigma^{2}$, we obtain
$\frac{d}{d(\sigma^{2})}f(\sigma^{2})$ $=\frac{1}{2\sigma^{2}}-\frac{\mathbb{E}_{X\sim p}\left[(X-\mu)^{2}\right]}{2(\sigma^{2})^{2}}$
$=\frac{1}{2(\sigma^{2})^{2}}\left(\sigma^{2}-\mathbb{E}_{X\sim p}\left[(X-\mu)^{2}\right]\right)$
Therefore, when $\sigma^{2}<\mathbb{E}_{X\sim p}\left[(X-\mu)^{2}\right]$, the derivative becomes negative. This implies that both $f(\sigma^{2})$ and consequently $KL(p||q)$ decrease as $\sigma^{2}$ increases. ∎
@@
[…]

### v2

@@ 4.1 Augment Concept Datasets by Modern Generative Models
In practice, it can be somewhat challenging to obtain a high‐quality, highly diverse dataset $\mathcal{C}_{\text{ori}}$ depicting the same concept, as required by our method. For example, as shown on the left side of Figure 2, the dataset provided by DreamBooth (Ruiz et al., 2023) contains four images of the same long‐eared corgi. Although the corgi is shown in various poses and from multiple viewpoints, the relatively uniform backgrounds do not provide sufficient diversity for our concept-based adversarial attack. Therefore, we decided to use Stable Diffusion XL (Podell et al., 2023) to expand the concept‐description dataset.
As illustrated in Figure 2, we designate the corgi as “[V] dog”. Using LoRA finetuning (Hu et al., 2022), we train an SDXL LoRA model on this concept. Next, we feed the five corgi images into GPT‐4o (Hurst et al., 2024), stating that these images represent the “[V] dog” and asking it to produce SDXL prompts that embody sufficient diversity for the “[V] dog.” Finally, we load the corgi LoRA into SDXL and, guided by GPT‐4o’s prompts (on the top of Figure 2), generate images featuring a wide range of viewpoints, environments, and poses for this corgi concept (see the right side of Figure 2).
@@ 4.2 Sample Selection
A key advantage of probabilistic adversarial attacks is that we can draw multiple samples from $p_{\text{adv}}$ and select the best ones [^3] Although deterministic methods may yield different results when sampled multiple times, their variability does not stem from algorithmic design but rather from other sources of error.. In a white-box scenario, we can simply discard samples that fail to deceive the classifier (rejection sampling). However, if $p_{\text{dis}}$ and $p_{\text{vic}}$ overlap only slightly, this may lead to high rejection rates (especially under a top-1 success criterion).
@@ 5.1 Data Preparation
We use the DreamBooth dataset (Ruiz et al., 2023), which provides 30 objects (animals, dolls, and everyday items) each with 5-6 representative images. To increase diversity, we apply the augmentation method in Section 4.1, generating 30 additional images per concept and forming the DreamBoothPlus dataset. Among the 30 objects of DreamBooth, we only augment 26, excluding four that pose challenges for text generation or require different fine-tuning parameters for cartoon-style content.
@@ 5.2 Fitting the Distance Distributions
We use the DreamBoothPlus dataset to finetune a diffusion model (Dhariwal and Nichol, 2021; Nichol and Dhariwal, 2021) to fit the distance distribution $p_{\text{dis}}$ (Details in Appendix D). We choose this model over more advanced architectures, such as the Stable Diffusion series (Podell et al., 2023; Rombach et al., 2022) or Flux, because it directly models $p(x)$ instead of $p(x\mid y)$, where $x$ is the image and $y$ is a label or prompt. Our goal is to employ a more principled model to illustrate our general adversarial attack method, rather than to optimize for the highest possible engineering performance.
@@ 5.4 Generating Targeted Adversarial Examples
We evaluate the performance of the concept-based adversarial attack in a targeted adversarial attack setting, because targeted attacks are generally more difficult than untargeted attacks [^4] This is especially true for ImageNet classifiers, which must distinguish among 1,000 classes. In an untargeted attack, the goal is simply to prevent the victim classifier from assigning the adversarial sample to its correct class. In contrast, a targeted attack requires the classifier to misclassify the adversarial sample exactly as the chosen target class $y_{\text{tar}}$.. In our experiments, we compare NCF (Yuan et al., 2022), ACA (Chen et al., 2024b), DiffAttack (Chen et al., 2024a), and ProbAttack (Zhang et al., 2024b). We include NCF because it is the strongest color-based adversarial attack. Both ACA and DiffAttack apply adversarial gradients in the latent space induced by Stable Diffusion and DDIM, representing the state of the art in unrestricted adversarial attacks. As discussed earlier in Section 3, ProbAttack can be viewed as a special case of our approach when $|\mathcal{C}_{\text{ori}}|=1$. For both ProbAttack and our approach, we set the number of samples $M$ to 10. During the sample selection phase of the experiments, our method uses two strategies — a conservative strategy and an aggressive strategy (both described in Section 4.2) — denoted in the tables as OURS (CONS) and OURS (AGGR), respectively.
@@
In our experiments, ResNet50 (He et al., 2016) is used as the victim classifier for white-box attacks. We measure transferability on VGG19 (Simonyan and Zisserman, 2015), ResNet152 (He et al., 2016), DenseNet161 (Huang et al., 2017), Inception V3 (Szegedy et al., 2016), EfficientNet B7 (Tan and Le, 2019), and on adversarially trained Inception V3 Adv (Kurakin et al., 2017), EfficientNet B7 Adv (Xie et al., 2020), and Ensemble IncRes V2 (Tramèr et al., 2018). We report both the white-box targeted attack success rate (on ResNet50) and the black-box transfer success rate (on the remaining models).
Table 1: Targeted attack success rates (%) on ImageNet classifiers. In the white-box setting, a targeted attack is counted as successful if the target class is ranked first. For transferability, we report top-5 success rates, counting an attack as successful if the target class is among the top 5 predictions (since top-1 success was uniformly low across all methods). See Appendix Q for full results.
@@
Transferability Targeted-Top5
VGG19 1.28 1.67 4.36 2.44 2.05 4.36
ResNet 152 1.41 1.92 8.33 3.33 2.82 8.72
DenseNet 161 1.41 2.05 7.44 3.97 3.85 11.54
Inception V3 0.90 1.41 3.08 2.56 1.28 4.74
EfficientNet B7 1.41 1.67 1.79 1.41 1.28 3.97
Adversarial Defence
Inception V3 Adv 1.15 1.28 3.21 2.18 0.90 3.72
EfficientNet B7 Adv 0.26 1.15 2.05 2.31 1.67 6.41
Ensemble IncRes V2 0.77 1.28 2.69 1.92 0.77 5.00
For unrestricted adversarial examples, we must also check whether they preserve the original concept and remain undetectable to humans. Therefore, we measure similarity via a user study (Appendix F) and CLIP (Radford et al., 2021), and image quality using no-reference metrics (MUSIQ (Ke et al., 2021), TReS (Golestaneh et al., 2022), NIMA (Talebi and Milanfar, 2018), ARNIQA (Agnolucci et al., 2024), DBCNN (Zhang et al., 2020), and HyperIQA (Su et al., 2020)).
As ImageNet has 1,000 classes, it is impractical to evaluate them all. Therefore, we randomly select 30 target classes $y_{\text{tar}}$, listed in Appendix E. Since DreamBoothPlus contains 26 concepts, each method generates $26\times 30=780$ adversarial examples. This scale is comparable to current popular approaches performing untargeted adversarial attacks on the ImageNet-Compatible dataset (Kurakin et al., 2018).
Table 1 presents the targeted attack success rates on ImageNet classifiers. Since the choice of aggressive or conservative strategy in the sample selection phase does not affect white-box performance, those rates are identical. Notably, the aggressive strategy achieves significantly higher transferability than other methods. While the conservative strategy leads to slightly lower transferability, it is still roughly comparable to the baseline methods. Please refer to Appendix Q for full results.
@@
$x_{\text{ori}}$ / $\mathcal{C}_{\text{ori}}$$y_{\text{tar}}$NCFACADiffAtkProbAtkOURS(CONS)OURS(AGGR)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionambulanceRefer to captionbasenjiRefer to captionhamsterRefer to captionking snakeRefer to captionpandaRefer to captiongoldfishRefer to captionmushroomRefer to captionlaptopRefer to captionpapillonRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 3: Qualitative comparison. (A green border indicates an example that successfully fools the classifier; red indicates failure.) See Appendix G for a more detailed qualitative analysis.
@@ 6 Related Works
To the best of our knowledge, no existing method constructs adversarial examples conditioned on an identity-level concept. The most closely related works are Song et al. (2018)’s, Dai et al. (2024)’s and Collins et al. (2025)’s. However, they treats a “class” (e.g., cat, dog, or truck) as the concept, which cannot precisely capture an individual identity. In our framework, we represent a concept through a distribution that could be learned from a set of images, allowing it to flexibly correspond to a single image (i.e. a set with size $1$), an identity-level concept, or a class-level concept. By contrast, other unrestricted adversarial attack methods focus solely on generating adversarial examples from a single image: Xiao et al. (2018a) generate adversarial examples using GANs, Chen et al. (2023) employ diffusion models, and Laidlaw et al. (2020) impose a feature-space distance as a regularization term in the optimization objective. ACA (Chen et al., 2024b) and DiffAttack (Chen et al., 2024a) further apply the attack gradient directly in the DDIM latent space (Song et al., 2020). Color-based transformations have proven effective in preserving semantic content for untargeted attacks (Bhattad et al., 2020; Hosseini and Poovendran, 2018; Shamsabadi et al., 2020; Yuan et al., 2022; Zhao et al., 2020), yet they perform poorly in targeted scenarios (Chen et al., 2024a), a result confirmed by our experiments.
Unrestricted Adversarial Attack Method Single-image Identity-level Concept Class-level Concept
Bhattad et al. (2020) Yes No No
Hosseini and Poovendran (2018) Yes No No
Colorfool (Shamsabadi et al., 2020) Yes No No
Zhao et al. (2020) Yes No No
NCF (Yuan et al., 2022) Yes No No
ACA (Chen et al., 2024b) Yes No No
DiffAttack (Chen et al., 2024a) Yes No No
ProbAttack (Zhang et al., 2024b) Yes No No
AdvGAN (Xiao et al., 2018a) Yes No No
Xiao et al. (2018b) Yes No No
Perceptual Adv. Attack (Laidlaw et al., 2020) Yes No No
Song et al. (2018) No No Yes
NatADiff (Collins et al., 2025) No No Yes
AdvDiff (Dai et al., 2024) No No Yes
AdvDiffuser Chen et al. (2023) Yes No Yes
Ours: Concept-based Adv. Attack Yes Yes Yes
Table 3: Comparison of unrestricted adversarial attack methods. Our method is the only one capable of performing adversarial attacks at the identity-level concept, while also supporting single-image and class-level concepts.
Our work directly inherited from Zhang et al. (2024b)’s probabilistic perspective, but we make a novel contribution by, for the first time, defining the distance distribution in adversarial attacks with respect to a distribution representing a concept rather than a single image. Although, operationally, our method appears to be a straightforward extension — replacing the single-image-centered distribution $p_{\text{dis}}(\cdot\mid x_{\text{ori}})$ with a concept-centered distribution $p_{\text{dis}}(\cdot\mid C_{\text{ori}})$ (as in the difference between (2) and (3)) — we rigorously demonstrate, both theoretically and empirically, why this seemingly simple generalization is remarkably effective.
@@ 7 Conclusions
Acknowledgments
This work was supported by the Engineering and Physical Sciences Research Council (EPSRC) through the AI Hub in Generative Models (EP/Y028805/1).
AZ and SK were funded by the UKRI AI Hub in Generative Models (EP/Y028805/1). SM was supported by the UKRI AI Programme Grant (EP/Y028856/1). AZ also acknowledges support from a personal grant provided by Mrs. Yanshu Wu. SK was additionally supported by the Research Council of Finland (Flagship Programme: Finnish Center for Artificial Intelligence FCAI, 359207) and the UKRI Turing AI World-Leading Researcher Fellowship (EP/W002973/1).
Ethics Statement
This work introduces a new class of adversarial attacks that operate at the concept level. While our primary goal is to advance scientific understanding of adversarial robustness and stimulate the development of stronger defenses, we acknowledge the potential for malicious misuse. In particular, concept-based adversarial attacks could be exploited to evade security-sensitive image classifiers or to manipulate systems deployed in safety-critical applications.
To mitigate these risks, we have:
  • Released all code and data strictly for research purposes, under licenses that encourage responsible use.
  • Discussed mitigation strategies in Appendix L, including adversarial training, AI-generated content detection, and hybrid defenses.
We emphasize that the broader impact (Appendix M) of this work depends on the research community’s response. By exposing vulnerabilities of current classifiers, we aim to encourage the development of more robust and trustworthy AI systems. We strongly discourage any use of this research for harmful purposes.
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our work:
  • Code and models: We provide the full source code, including scripts for dataset preparation, model fine-tuning, and adversarial example generation, at https://github.com/andiac/ConceptAdv. All hyperparameters and training details are specified in the code repository.
  • Datasets: Our experiments are based on the DreamBooth dataset (Ruiz et al., 2023), which is publicly available under a CC-BY-4.0 license. We also describe our augmentation procedure using SDXL and LoRA in Section 4.1, and provide scripts for generating data (DreamBoothPlus) as part of our code repo.
  • Hyperparameters: Fine-tuning settings for both SDXL LoRA and diffusion models are detailed in Appendix D. For sampling and evaluation, we report the number of generated adversarial examples, sampling strategies, and evaluation metrics in Sections 4-5.
  • Theoretical results: Proofs of all theorems are included in Appendix A, and additional details on KL divergence estimation are in Appendix B.
  • Compute resources: We report hardware specifications and training times in Appendix I to allow others to reproduce our experiments with similar resources.
We believe these resources provide sufficient detail for reproducing both our theoretical and empirical results.
Appendix
@@ Appendix A Proof of the Theorems
Let $p$ be a probability distribution and $q$ be a Gibbs distribution of the form
$q(x)=\frac{\exp(-\beta D(x,\mu))}{Z(\beta)},$
where $Z(\beta)$ is the normalizing constant, $\mu$ is a constant and $D$ is a distance function. Then $KL(p\,\|\,q)$ is a increasing function of $\beta$ whenever $\mathbb{E}_{X\sim p}[D(X,\mu)]>\mathbb{E}_{X\sim q}[D(X,\mu)]$.
@@
Since $\int p(x)\log p(x)dx$ is independent of $\beta$, we can treat it as a constant. Let us denote the $\beta$-dependent component as $f(\beta)$, which gives us
$f(\beta)$ $=-\int p(x)\log q(x)dx$
$=-\int p(x)\left[-\beta D(x,\mu)-\log Z(\beta)\right]dx$
$=\int p(x)\beta D(x,\mu)dx+\log Z(\beta)\int p(x)dx$
$=\mathbb{E}_{X\sim p}[\beta D(X,\mu)]+\log Z(\beta)$
Taking the derivative with respect to $\beta$, we obtain
$\frac{d}{d\beta}f(\beta)$ $=\mathbb{E}_{X\sim p}[D(X,\mu)]+\frac{1}{Z(\beta)}\frac{dZ(\beta)}{d\beta}$
$=\mathbb{E}_{X\sim p}[D(X,\mu)]+\frac{1}{Z(\beta)}\int\frac{d\exp(-\beta D(X,\mu))}{d\beta}dx$
$=\mathbb{E}_{X\sim p}[D(X,\mu)]+\frac{1}{Z(\beta)}\int-D(X,\mu)\exp(-\beta D(X,\mu))dx$
$=\mathbb{E}_{X\sim p}[D(X,\mu)]-\mathbb{E}_{X\sim q}[D(X,\mu)]$
Therefore, when $\mathbb{E}_{X\sim p}[D(X,\mu)]>\mathbb{E}_{X\sim q}[D(X,\mu)]$, the derivative becomes positive. This implies that both $f(\beta)$ and consequently $KL(p||q)$ increase as $\beta$ increases. ∎
@@
$p_{\text{vic}}(x\mid y_{\text{tar}})\;=\;\frac{\exp(-c\,f(x,y_{\text{tar}}))}{\int\exp(-c\,f(x,y_{\text{tar}}))\,dx}.$
Let $Z:=\int\exp\!\bigl(-c\,f(x,y_{\text{tar}})\bigr)\,dx$. Then